OpenStudy (anonymous):

How did I do? Use a power series to approximate the definite integral to six decimal places. \[\int_{0}^{0.2}\frac{1}{1+x^5}dx\] \[\int_{0}^{0.2}\frac{1}{1-(-x^5}dx\] \[\int_{0}^{0.2}\sum_{n=0}^{\infty}(-x^5)^n dx\] \[\int_{0}^{0.2}\sum_{n=0}^{\infty}(-1)^nx^{5n} dx\] \[\sum_{n=0}^{\infty}\left[(-1)^n\frac{x^{5n+1}}{5n+1} \right]_0^{0.2}\]

OpenStudy (anonymous):

\[ \Large \frac{1}{1-x} = \sum_{n=0}^\infty x^n\] You used this right?

OpenStudy (anonymous):

Seems pretty correct to me.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I have a dumb question, so do I sub those limits in for n or x?

OpenStudy (anonymous):

x of course because the limit is from x=0 to x=0.2... so what would I do with the n?

OpenStudy (anonymous):

yes, but there is your problem, what you did here is correct. But usually you write the terms out, understand what I mean? So if you want to make a rough approximation you have to carry out the steps.

OpenStudy (anonymous):

I mean the integral is from 0 to 0.2

OpenStudy (anonymous):

Let me try to write it down, I better post a cheat sheet before I make some careless mistakes.

OpenStudy (anonymous):

so when n=0 and n=1 and n=2 and so on and so forth?

OpenStudy (anonymous):

http://mathworld.wolfram.com/MaclaurinSeries.html

OpenStudy (anonymous):

I did post that for myself too by the way (-:

OpenStudy (anonymous):

but it's power series isn't it?

OpenStudy (anonymous):

Sooo \[ \Large \frac{1}{1-x}=1+x+x^2+x^3+x^4+x^5 + \cdots \] The series is correct you have choosen, but you have found the most general answer, and as far as I know, I can't work with that to find a approximate solution. My idea: Make the same substitution for the Series above (not the general form) and then evaluate each term individually by taking the integral.

OpenStudy (anonymous):

I feel soo unintelligent right now...why 1+x+x^2 etc.?

OpenStudy (anonymous):

well that is the series. http://mathworld.wolfram.com/MaclaurinSeries.html I hope I didn't choose the wrong one.

OpenStudy (anonymous):

\[ \Large \frac{1}{1-x}=1+x+x^2+x^3+ \cdots = \sum_{n=0}^\infty x^n \]

OpenStudy (anonymous):

See, one is exact, right hand side, one is approximate, left hand side (middle)

OpenStudy (anonymous):

So if you want to approximate, you don't want to use the 'exact' general case. I hope that makes some sense.

OpenStudy (anonymous):

oh it makes sense now

OpenStudy (experimentx):

put the values of x alternating series can be approximated this way http://en.wikipedia.org/wiki/Alternating_series#Approximating_sums

OpenStudy (anonymous):

nevertheless @MathSofiya , just so you know, what you did above is not for nothing, you can still see how your new series looks like by running it through the way you have defined it, in other words substitute a few of the n terms like n=0, n=1 ... and then you can integrate that term by term

OpenStudy (anonymous):

how far should I go? till n=5 maybe?

OpenStudy (anonymous):

Erm usually I do this via try and error, maybe there is a more elegant way *grins* let me see real quick before I say something that I will later have to pull back heh.

OpenStudy (anonymous):

I would say just take a few of them, and then when you evaluate your integral at zero, all terms will go to zero, so evaluate each of it at higher bound and then see how many decimals each successive integral will give.

OpenStudy (experimentx):

use this ... the tn must be of 6th order ... all (converging) alternating series can be approximated this way.

OpenStudy (anonymous):

ok I'll try it and I'll post what i get

OpenStudy (anonymous):

Man this is difficult :\ \[(-1)^n\frac{(0.2)^{5n+1}}{5n+1}-(-1)^n\frac{(0)^{5n+1}}{5n+1}\] and now when I do the series I use \[\left[(-1)^3\frac{(0.2)^{5*3+1}}{5*3+1}+(-1)^2\frac{(0.2)^{5*2+1}}{5*2+1}+(-1)^1\frac{(0.2)^{5*1+1}}{5*1+1}+(-1)^0\frac{(0.2)^{5*0+1}}{5*0+1}\right]\]- \[\left[(-1)^3\frac{(0)^{5*3+1}}{5*3+1}+(-1)^2\frac{(0)^{5*2+1}}{5*2+1}+(-1)^1\frac{(0)^{5*1+1}}{5*1+1}+(-1)^0\frac{(0)^{5*0+1}}{5*0+1}\right]\]

OpenStudy (anonymous):

something like that?

OpenStudy (anonymous):

well, it looks correct to me at first glance @MathSofiya , although you could have saved the second line (the subtraction) because it's all evaluated at x=0, therefore all the terms become zero.

OpenStudy (anonymous):

I don't quite understand what you mean by it's evaluated at x=0

OpenStudy (anonymous):

By the way, this is the exact result http://www.wolframalpha.com/input/?i=integrate+1%2F%281%2Bx%5E5%29 , you can click on "Show steps", it's not less painful.

OpenStudy (anonymous):

well, taylor/maclaurin series are polynomial approximations, polynomials without a constant c evaluated at zero are zero.

OpenStudy (anonymous):

am i totally wrong, or for the first one would you not simply use \(1-x^5\) ?

OpenStudy (anonymous):

where @satellite73 ?

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

Like, lets say you have integrated some ominous function that I don't want to define any more carefully now, for a definite integral and you have obtained the following result: \[ \Large F(x)= \left. x^7+3x^3+2x+\right|_0^3 \] If you plug in zero all the terms with an x will vanish.

OpenStudy (anonymous):

you mean the second line...all the way at the top of the page?

OpenStudy (anonymous):

oh I see @spacelimbus. So since we don't have any constants, whatever is to the right of the minus sign will disappear.

OpenStudy (anonymous):

yeah i came here late but \(\frac{1}{1+x^5}=1-x^5+x^{10}-...\) expanded at 0, and you are only going out to 6 decimal places, which means you can ignore the \((.2)^{10}\) part

OpenStudy (anonymous):

yes, the minus will disappear.

OpenStudy (anonymous):

how did you come to the conclusion that (.2)^10 is the "limit"

OpenStudy (anonymous):

or i mean that's where i would stop

OpenStudy (anonymous):

so i think (correct me if i am wrong) you just need to evaluate \(x-\frac{x^6}{6}\) at \(x=.2\)

OpenStudy (anonymous):

shoot me. Really?

OpenStudy (anonymous):

you are only looking for 6 decimal place accuracy right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\((.2)^{11}\) is some little tiny number beyond the sixth decimal place

OpenStudy (anonymous):

oh. I see... so back to the \[x-\frac{x^6}{6}\] how and why?

OpenStudy (anonymous):

i didn't mean to butt in so late and have not really read all the above posts, but the first one looked pretty straightforward

OpenStudy (anonymous):

\(\frac{1}{1+x^5}\) expands as \(1-x^5+x^{10}-x^{15}+...\)

OpenStudy (anonymous):

No problem @satellite73 , I believe the derivation we have done is correct, for the approximation of the given integral, I just don't know how to "bind" the problem to given decimals.

OpenStudy (anonymous):

that is, take well known expansion for \(\frac{1}{1-x}\) and replace \(x\) by \(-x^5\)

OpenStudy (anonymous):

now integrate term by term

OpenStudy (anonymous):

for whatever reason \[1-x^5+x^{10}-x^{15}+...\] is still a magical occurrence

OpenStudy (anonymous):

you get \[x-\frac{x^6}{6}+\frac{x^{11}}{11}-...\]

OpenStudy (anonymous):

I can do the integral, but the first step is still a mystery

OpenStudy (anonymous):

I believe it's exactly what we derived isn't it? Sorry if I increase the confusion,

OpenStudy (anonymous):

and since you only are looking for six decimal place accuracy, you can ignore everything beyond \(\frac{x^6}{6}\)

OpenStudy (anonymous):

the gimmick is always to start with \[\frac{1}{1-x}=\sum_{k=0}^{\infty}x^k\]

OpenStudy (anonymous):

make whatever you have look like that if you can

OpenStudy (anonymous):

I did that and came up with \[\sum_{n=0}^{\infty}(-1)^nx^{5n} dx\]

OpenStudy (anonymous):

without the dx of course

OpenStudy (anonymous):

oooohhhh. I think i see it now

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

oh you wrote them out before you integrated...

OpenStudy (anonymous):

\[1-x^5+x^{10}-x^{15}-...\]

OpenStudy (anonymous):

then integrate term by term

OpenStudy (anonymous):

and evaluate at x=0.2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and notice that \(\frac{.2^{11}}{11}\) is way beyond the sixth decimal place

OpenStudy (anonymous):

but the way that @Spacelimbus and i did is correct too?

OpenStudy (anonymous):

i believe so, yes

OpenStudy (anonymous):

in fact, what @satellite73 wrote here, I have mentioned above too, I believe we just have had some misunderstanding. *repost* yes, but there is your problem, what you did here is correct. But usually you write the terms out, understand what I mean? So if you want to make a rough approximation you have to carry out the steps.

OpenStudy (anonymous):

Mystery solved guys! Thank you so much Everyone!

OpenStudy (anonymous):

I wasn't strict enough I believe which caused some confusion (-: So thanks for cleaning things up @satellite73

OpenStudy (anonymous):

<---can kitten have medal too?

OpenStudy (anonymous):

sure, but in the next post because i used mine up

OpenStudy (anonymous):

lol that's ok :P

OpenStudy (anonymous):

I will have to safe mine too it's already gone here (=

OpenStudy (anonymous):

no prob. Thanks for helping